# Discrete Diffusion in Large Language and Multimodal Models: A Survey

### v1

@@ V-C2 KV-Cache
The general experimental observations are that caching in dLLMs leads to performance fluctuations, and in most cases, results in performance degradation. The severity of degeneration increases with larger update intervals. However, with smaller update intervals (e.g., 2 to 8), performance degradation is minimal and often comparable to that of models without caching.
V-D Guidance Techniques
@@
V-D1 Classifier-Free Guidance
@@ VI Applications
VI-A Text Generation and Style Control
@@
VI-B Text Editing and Summarization
@@
VI-C Sentiment Analysis and Data Augmentation
@@
The $\textrm{CDA}^{2}$ framework [55] employs counterfactual diffusion augmentation to improve cross-domain adaptation in low-resource sentiment analysis. It generates counterfactual target samples by substituting domain-relevant words in source data and guiding the diffusion model to produce high-quality target domain examples. Experiments show that $\textrm{CDA}^{2}$ generates higher-quality target samples and achieves state-of-the-art performance in cross-domain sentiment tasks. DiffusionCLS [56] examines the use of a diffusion language model for data augmentation in low-resource sentiment classification. It generates pseudo-samples by reconstructing strongly sentiment-related tokens, thereby increasing data diversity while preserving label-consistent information. This approach balances consistency and diversity without introducing excessive noise. Pushing diffusion further into structured sentiment tasks, the Grid Noise Diffusion Pinpoint Network (GDP) [57] introduces a novel diffusion-based and template-guided approach for Aspect Sentiment Quad Prediction. It includes three modules: Diffusion Vague Learning to enhance learning under uncertainty, Consistency Likelihood Learning to capture commonalities of sentiment elements and mitigate noise, and a GDP-FOR generation template for more natural outputs. Extensive experiments demonstrate the remarkable effectiveness of GDP on ASQP tasks.
VI-D Knowledge and Reasoning
@@
Reasoning is one of the most important abilities of models in the large model era. In the dLLM scenario, some related works also emerged. Specifically, Diffusion-of-Thought (DoT) [58] firstly integrates chain-of-thought reasoning into dLLMs to enhance reasoning capabilities. Unlike traditional left-to-right autoregressive reasoning, DoT allows reasoning steps to diffuse over multiple steps, offering a flexible trade-off between compute and reasoning performance. In addition, DiffuCOMET [59] develops a series of models that leverage diffusion to infer contextually relevant commonsense knowledge from narratives. The method progressively refines a commonsense fact representation anchored to the input context through multiple diffusion steps, producing inferences that are both contextually appropriate and diverse. On benchmarks like ComFact and WebNLG+, DiffuCOMET achieves a better trade-off between commonsense diversity, contextual relevance, and alignment to known facts. DPCL-Diff [60] combines graph node diffusion with dual-domain periodic contrastive learning for temporal knowledge graph reasoning. Its Graph Node Diffusion (GNDiff) model introduces noise into sparsely related historical events to generate new event data that better matches the true distribution, enhancing the ability to reason about future events; the dual-domain learning maps periodic event entities to a Poincaré space and non-periodic entities to a Euclidean space to improve representation power. Most recently, the d1 framework [61] adapts pretrained dLLMs into reasoning models via a combination of supervised fine-tuning and reinforcement learning. It introduces a novel critic-free policy-gradient algorithm called diffu-GRPO and employs masked SFT to distill reasoning knowledge directly from existing datasets, thereby enhancing the model’s reasoning abilities.
VI-E Vision and Multimodal
@@
DiffVLA [62] introduces a vision-language-guided diffusion policy for autonomous-driving planning. Specifically, it employs a hybrid sparse–dense diffusion strategy, guided by a vision-language model, thereby improving both efficiency and action diversity in multi-modal driving scenarios. In a image-enhancement context, UDAN-CLIP [63] proposes an image-to-image diffusion framework for underwater enhancement that jointly integrates a vision-language model and a spatial-attention module. During pre-training, the model leverages synthetic underwater data and a CLIP-guided loss to retain in-air natural priors while correcting local degradations such as haze and low contrast. Extending diffusion policies to robotics, [64] present a framework VPDD that first pre-trains on large-scale actionless human videos and then transfers the learned discrete diffusion policy to robot-control tasks. Here, both human and robot videos are encoded into unified video tokens; a discrete diffusion model is trained generatively on the human corpus before being fine-tuned on a small, action-labeled robot set. Turning to motion synthesis, M2D2M [65] employs a discrete diffusion model to generate continuous human-motion sequences from textual descriptions of multiple actions. By adapting dynamic transition probabilities to token-level similarities and adopting a two-phase sampling strategy—independent followed by joint denoising—it produces long-term, smooth, and contextually coherent motions, ultimately outperforming state-of-the-art multi-motion baselines. Finally, AR-Diffusion [66] introduces a novel architecture that blends autoregressive and diffusion techniques for flexible, asynchronous video generation. It gradually corrupts video frames during both training and inference to reduce phase discrepancies; meanwhile, a non-decreasing length-control mechanism—borrowed from autoregressive generation—enables consistent handling of variable-length sequences, thereby enhancing temporal coherence.
VI-F Biological and Drug Discovery

### v2

@@ V.C.2 KV-Cache
  • DualCache [51]. DualCache adopts a block-wise caching strategy, decoding text block by block. The block currently being decoded is not cached, while the surrounding blocks are cached.
@@
The general experimental observations are that caching in dLLMs leads to performance fluctuations, and in most cases, results in performance degradation. The severity of degeneration increases with larger update intervals. However, with smaller update intervals (e.g., 2 to 8), performance degradation is minimal and often comparable to that of models without caching [25, 26].
V.D Guidance Techniques
@@
V.D.1 Classifier-Free Guidance
@@ V.D.3 Reward Guidance
V.D.4 Energy-Based Diffusion
xu2025energybased [53] proposes the Energy-based Diffusion Language Model (EDLM), which augments a pretrained diffusion model $p_{\theta}(x_{0}\mid x_{t})$ with an unnormalized energy model $E_{\phi}(x_{0},x_{t},t)$, yielding a joint denoising distribution:
$p_{\theta,\phi}(x_{0}\mid x_{t})=p_{\theta}(x_{0}\mid x_{t})\cdot\frac{\exp\left(-E_{\phi}(x_{0},x_{t},t)\right)}{Z_{\phi}(x_{t})},$ (97)
where $Z_{\phi}(x_{t})$ is the intractable partition function required for normalization. This residual formulation corrects the denoising distribution by reweighting samples from the diffusion model using the energy function. The energy function can be derived from either a pretrained autoregressive (AR) model or a finetuned diffusion model. To implement this energy function-based guidance, [53] adopts an importance sampling strategy. The decoding process at each time step can be summarized as follows.
  1. 1.
    Generate $k$ candidate samples $\{x_{0}^{(i)}\}_{i=1}^{k}\sim p_{\theta}(x_{0}\mid x_{t})$ using the diffusion model.
  2. 2.
    Compute the unnormalized energy scores $e^{(i)}=E_{\phi}(x_{0}^{(i)},x_{t},t)$ for all samples in $\{x_{0}^{(i)}\}_{i=1}^{k}$.
  3. 3.
    Sample one $x_{0}$ from the candidate pool according to importance weights:
    $w^{(i)}=\frac{\exp(-e^{(i)})}{\sum_{j=1}^{k}\exp(-e^{(j)})}.$ (98)
  4. 4.
    Use the sampled $x_{0}$ to perform one denoising step via the backward posterior:
    $x_{t-1}\sim q(x_{t-1}\mid x_{t},x_{0}).$ (99)
@@ VI Applications
VI.A Text Generation and Style Control
@@
VI.B Text Editing and Summarization
@@
VI.C Sentiment Analysis and Data Augmentation
@@
The $\textrm{CDA}^{2}$ framework [63] employs counterfactual diffusion augmentation to improve cross-domain adaptation in low-resource sentiment analysis. It generates counterfactual target samples by substituting domain-relevant words in source data and guiding the diffusion model to produce high-quality target domain examples. Experiments show that $\textrm{CDA}^{2}$ generates higher-quality target samples and achieves state-of-the-art performance in cross-domain sentiment tasks. DiffusionCLS [64] examines the use of a diffusion language model for data augmentation in low-resource sentiment classification. It generates pseudo-samples by reconstructing strongly sentiment-related tokens, thereby increasing data diversity while preserving label-consistent information. This approach balances consistency and diversity without introducing excessive noise. Pushing diffusion further into structured sentiment tasks, the Grid Noise Diffusion Pinpoint Network (GDP) [65] introduces a novel diffusion-based and template-guided approach for Aspect Sentiment Quad Prediction. It includes three modules: Diffusion Vague Learning to enhance learning under uncertainty, Consistency Likelihood Learning to capture commonalities of sentiment elements and mitigate noise, and a GDP-FOR generation template for more natural outputs. Extensive experiments demonstrate the remarkable effectiveness of GDP on ASQP tasks. [106] addresses the layout sticking issue in discrete diffusion layout generation, where poorly placed elements become hard to fix. It proposes Layout-Corrector, a simple module that scores elements’ layout quality, resets poorly positioned tokens, and guides refined placement conditioned on high-quality elements. Combined with discrete diffusion models, Layout-Corrector consistently enhances layout quality, flexibility, and sampling efficiency.
VI.D Knowledge and Reasoning
@@
Diffusion-of-Thought (DoT) [66] firstly integrates chain-of-thought reasoning into dLLMs to enhance reasoning capabilities. Unlike traditional left-to-right autoregressive reasoning, DoT allows reasoning steps to diffuse over multiple steps, offering a flexible trade-off between compute and reasoning performance. In addition, DiffuCOMET [67] develops a series of models that leverage diffusion to infer contextually relevant commonsense knowledge from narratives. The method progressively refines a commonsense fact representation anchored to the input context through multiple diffusion steps, producing inferences that are both contextually appropriate and diverse. On benchmarks like ComFact and WebNLG+, DiffuCOMET achieves a better trade-off between commonsense diversity, contextual relevance, and alignment to known facts. DPCL-Diff [68] combines graph node diffusion with dual-domain periodic contrastive learning for temporal knowledge graph reasoning. Its Graph Node Diffusion (GNDiff) model introduces noise into sparsely related historical events to generate new event data that better matches the true distribution, enhancing the ability to reason about future events; the dual-domain learning maps periodic event entities to a Poincaré space and non-periodic entities to a Euclidean space to improve representation power. Most recently, the d1 framework [69] adapts pretrained dLLMs into reasoning models via a combination of supervised fine-tuning and reinforcement learning. It introduces a novel critic-free policy-gradient algorithm called diffu-GRPO and employs masked SFT to distill reasoning knowledge directly from existing datasets, thereby enhancing the model’s reasoning abilities. [49] provides an insight into DLLM reasoning: the difficulty of decoding individual tokens in a response varies, and it is not necessarily the case that tokens on the left are easier to decode than those on the right. Therefore, autoregressive (AR) reasoning may not be the optimal approach. Diffusion-based LLMs demonstrate clear advantages over AR models in tasks involving planning, reverse reasoning, and global context understanding. [49] presents examples such as Sudoku and the 24 Game to illustrate these points. NeSyDM [107] introduces a discrete diffusion process in the symbolic space to overcome the limitations of traditional neurosymbolic systems, which often rely on the assumption of conditional independence between symbols. The framework first extracts symbolic representations from perceptual inputs using a neural network, and then performs a multi-step denoising diffusion process over these symbols. This allows the model to explicitly capture inter-symbol dependencies through the structure of the diffusion trajectory.
VI.E Vision and Multimodal
@@
DiffVLA [70] introduces a vision-language-guided diffusion policy for autonomous-driving planning. Specifically, it employs a hybrid sparse–dense diffusion strategy, guided by a vision-language model, thereby improving both efficiency and action diversity in multi-modal driving scenarios. In a image-enhancement context, UDAN-CLIP [71] proposes an image-to-image diffusion framework for underwater enhancement that jointly integrates a vision-language model and a spatial-attention module. During pre-training, the model leverages synthetic underwater data and a CLIP-guided loss to retain in-air natural priors while correcting local degradations such as haze and low contrast. Extending diffusion policies to robotics, [72] present a framework VPDD that first pre-trains on large-scale actionless human videos and then transfers the learned discrete diffusion policy to robot-control tasks. Here, both human and robot videos are encoded into unified video tokens; a discrete diffusion model is trained generatively on the human corpus before being fine-tuned on a small, action-labeled robot set. Turning to motion synthesis, M2D2M [73] employs a discrete diffusion model to generate continuous human-motion sequences from textual descriptions of multiple actions. By adapting dynamic transition probabilities to token-level similarities and adopting a two-phase sampling strategy—independent followed by joint denoising—it produces long-term, smooth, and contextually coherent motions, ultimately outperforming state-of-the-art multi-motion baselines. Finally, AR-Diffusion [74] introduces a novel architecture that blends autoregressive and diffusion techniques for flexible, asynchronous video generation. It gradually corrupts video frames during both training and inference to reduce phase discrepancies; meanwhile, a non-decreasing length-control mechanism—borrowed from autoregressive generation—enables consistent handling of variable-length sequences, thereby enhancing temporal coherence. Besides, the understanding tasks, discrete diffusion is also largely used in the vision generation tasks [47, 103].
VI.F Biological and Drug Discovery